Brown Cow Dairy uses the aging approach to estimate Bad Debt Expense. The balance of each account receivable is aged on the bWhat amount of Bad Debt Expense should be recorded on December 31? Amount of Bad Debt ExpenseRequired 2 Required 3 Required 1 If the unadjusted balance in the Allowance for Doubtful Accounts was a $590 debit balance, w
Answer #1
1-30 31-90 >90 Total
Account receivable 11900 4900 2900 19700
Estimated uncollectible 4% 12% 15%
Estimated uncollectible 476 588 435 1499

2) Bad debt expense = 1499-790 = 709

3) Bad debt expense = 1499+590 = 2089